OpenStudy (anonymous):

Find the midpoint of the segment with the given endpoints. D(–5, 2) and E(–13, 0)

OpenStudy (anonymous):

do you need the coordinates of the midpoint?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the formula for midpoint is \[\left( \frac{ x _{1}+x _{2} }{2 }, \frac{ y _{1}+y _{2} }{ 2 } \right)\] just plug in your given points

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

np:)